### v1

# Anomalous nuclear spin coherence in superconducting Nb3Sn

Abstract
We have investigated the normal and superconducting states of the technologically important compound Nb3Sn using 93Nb nuclear magnetic resonance. From spin-lattice relaxation we find strong suppression of the zero-temperature superconducting order parameter by magnetic field. We have identified an anomalously large electron-nuclear exchange interaction from spin-spin relaxation measurements, an order of magnitude beyond that of the dipole-dipole interaction, and thereby sensitive to vortex dynamics and vortex pinning.
I Introduction
Nb3Sn is a superconductor with high critical temperature $T_{c}(\sim 18\,$K) and high critical field $H_{c2}(\sim 30\,$T). [1, 2, 3] Its unique physical properties have led to applications in superconducting magnet technology, especially for achieving very high magnetic fields. Little is known about this compound from the microscopic perspective provided by nuclear magnetic resonance, 93Nb (NMR). In our previous brief work, we surveyed the 93Nb NMR properties of Nb3Sn in both normal and superconducting states. [4] Here we report detailed measurements of the spin-lattice, i.e. longitudinal, relaxation rate ($T_{1}^{-1}$) at low temperatures to provide a measure of the average amplitude of the superconducting order parameter and its suppression by magnetic field. However, the focus in the present work, is the temperature dependence of the spin-spin, i.e. transverse, relaxation rate ($T_{2}^{-1}$). Previous NMR studies of superconducting compounds, most notably in high-$T_{c}$ superconductors,[5, 6] has related $T_{2}$ to vortex dynamics. This is also possible for Nb3Sn owing to its anomalously high nuclear spin coherence.
The measure of nuclear spin coherence is expresssed by the relaxation time $T_{2}$ which is determined by magnetic field fluctuations at relatively low frequencies, of order kHz, in the direction of the steady magnetic field. In metallic materials magnetic field fluctuations can couple to the nuclear spin through the hyperfine interaction via conduction electrons. In special cases this Ruderman-Kittel indirect exchange interaction (RKKY)[7] between like nuclear spins reduces the nuclear dipole-dipole interaction leading to substantial narrowing of the NMR spectrum; often described as being similar to extreme motional narrowing in liquids.[8, 9] The prime example in a metal is 195Pt NMR in pure platinum.[10] We discovered that Nb3Sn is in the same class of such systems with an unusually long $T_{2}$, perhaps not surprising since niobium has a single, non-zero spin, isotope 93Nb ($I=9/2$), a requirement for this phenomenon. We discuss this later, and, most importantly, the effect of the superconducting state on the RKKY interaction.
II Sample preparation and characterization
The high-quality powder sample was produced in the Applied Superconductivity Center of the National High Magnetic Field Laboratory following a process of high-energy ball milling, initial cold isostatic press densification, and final hot isostatic press reaction and densification [2]. In this work we have two samples with slightly different preparation process. The parameters of the two samples are listed in Table  1. According to X-ray diffraction (XRD) characterization, Nb3Sn has a typical cubic A15 crystal structure, shown in Fig. 1. The Sn atoms sit on the corners and the body-centered positions of a unit cell, while the Nb atoms occupy three orthogonal chains on the faces of the cube.
Figure 1: Crystal structure of stoichiometric Nb3Sn.
Sample Composition Heat Treatment $H_{c2}(0)$ $T_{c}$ at 0 field Weight
No.1 27% Sn 1800 °C 28.5 T 17.4 K 20 mg
No.2 25% Sn 1600 °C 30.7 T 18.0 K 220 mg
Table 1: A summary of Nb3Sn powder samples used for NMR measurements; also Supplementary Materials.
III Experiment
The NMR experiments were performed at the National High Magnetic Field Laboratory with magnetic fields varying from 3.7 T to 15 T. The superconducting transition temperature $T_{c}$ at different fields was identified from heat capacity measurement, which is consistent with the detuning of the NMR coil at onset of superconductivity.
The $T_{2}^{-1}$ was measured by a $90^{\circ}-180^{\circ}$ pulse sequence, known as the Hahn echo [11], with varying time interval $\tau$ between the two pulses. The $T_{2}^{-1}$ was obtained by fitting the recovery curve, which will be discussed later. The $T_{1}^{-1}$ is measured using a $90^{\circ}-\tau-90^{\circ}-\tau_{0}-180^{\circ}$ pulse sequence, where the Hahn echo is acquired with a fixed $\tau_{0}$ from a pure Lorentzian recovery.
IV Results
IV.1 Longitudinal Relaxation
The longitudinal relaxation rate $T_{1}^{-1}$ was measured at the central transition in the superconducting state at 3.7 T, 7.5 T, 11.3 T and 15.0 T respectively in Sample 1. It was found to be the same at satellite frequencies. The $T_{1}^{-1}$ was obtained by fitting the recovery curve $M(t)$ to a stretched ”Master equation” [12] with $\beta\sim 0.9$ for the five pairs of quadrupolar satellites:
$\begin{split}M(t)=M_{0}-&M_{0}[\frac{1}{165}\mathrm{exp}(-t/T_{1})^{\beta}+\frac{24}{715}\mathrm{exp}(-6t/T_{1})^{\beta}\\ &+\frac{6}{65}\mathrm{exp}(-15t/T_{1})^{\beta}+\frac{1568}{7293}\mathrm{exp}(-28t/T_{1})^{\beta}\\ &+\frac{7938}{12155}\mathrm{exp}(-45t/T_{1})^{\beta}]\end{split}$ (1)
Figure 2: 93Nb longitudinal relaxation rate $T_{1}^{-1}$ in the superconducting state of the Nb3Sn powder (Sample 1.). The lines are the fitting results using Eq. 2 and Eq. 3 at different magnetic fields. The slope approximation in the low temperature region represents the average energy gap at zero temperature ($\Delta(0)$). Inset: $\Delta(0)/k_{B}T_{c}$ vs. external field where $\Delta(0)$ comes from the fitting, the dashed line is a guide to the eye.
The temperature dependence of $T_{1}^{-1}$ in the low temperature region is plotted in Fig. 2. The data are fitted by Eq. 2, identifying the dependence of $T_{1}^{-1}$ on the superconducting energy gap $\Delta(T)$, and Eq. 3, illustrating the temperature dependence of the gap with strong-coupling taken into account [13]:
$T_{1}^{-1}\propto\mathrm{exp}(-\Delta(T)/k_{B}T)$ (2)
$\Delta(T)=\Delta(0)\mathrm{tanh}\{\frac{\pi k_{B}T_{c}}{\Delta(0)}[\frac{2}{3}(\frac{T_{c}}{T}-1)\frac{\Delta C}{C}]^{1/2}\},$ (3)
where the heat capacity jump $\Delta C/C$ incorporates strong-coupling in a BCS energy gap.
The energy gap in the zero temperature limit $\Delta(0)$ was obtained, taking the temperature dependence of $T_{1}^{-1}$ and $\Delta C/C$ from heat capacity measurements. This is plotted in the inset of Fig. 2 as a function of $T_{c}/T$. We observe a clear decrease in the energy gap as a function of the external magnetic field, indicating suppression of the order parameter by field. This suppression behavior is also seen in V3Si from STM measurements.[14] The $\Delta(0)$ at 15 T is even lower than the BCS weak coupling value $1.76\,k_{B}T_{c}$. In this temperature range there is no evidence of a linear temperature term associated with vortex cores.[15]
IV.2 Transverse Relaxation
Figure 3: Temperature dependence of 93Nb transverse relaxation rate $T_{2}^{-1}$ of the Nb3Sn powder in various fields (Sample 1.). (a) Relaxation rate $T_{2}^{-1}$ vs. Temperature. The lines are guides to the eye. Inset: $T_{2}^{-1}$ Jump at $T_{c}$ in different fields. (b) The same relaxation rate $T_{2}^{-1}$ vs. $T/T_{c}(H)$. Plotting versus temperature and magnetic field clearly shows common aspects in both the superconducting and normal states in this sample. The dashed lines in the normal state are linear fits and in the superconducting state a guide to the eye. The extrapolations to $T=0$ are notable.
The transverse relaxation rate $T_{2}^{-1}$ was measured at the central transition in both the normal and superconducting states at various fields. For Sample 1, this is shown in Fig. 3. The $T_{2}^{-1}$ was obtained from a stretched exponential fit to measured relaxation curves. In Fig. 3(a,b), $T_{2}^{-1}$ exhibits a linear dependence on temperature in the normal state, despite the fact that the nuclear dipole-dipole interaction is temperature-independent. However, in addition to the dipole-dipole interaction, $T_{2}^{-1}$ is also constrained by the lifetime of the excited nuclear spin, referred to as the ”Redfield effect” [16, 9], which explicitly takes into account the dependence of $T_{2}^{-1}$ on $T_{1}^{-1}$. From the Korringa law, $T_{1}^{-1}$ is linear in the normal state, consequently this linear behavior of $T_{2meas}^{-1}$ strongly suggests a Redfield effect expressed as,
$T_{2meas}^{-1}=T_{2}^{-1}+\kappa\cdot T_{1i}^{-1}$ (4)
where $T_{2meas}^{-1}$ is the measured echo response and $\kappa\cdot T_{1i}^{-1}$ represents the Redfield term. The rate $T_{1i}^{-1}$ is the initial component of the energy relaxation over the same time scale as $T_{2}^{-1}$, corresponding to the first-order approximation in Eq. 1 since $T_{2}\,<<\,T_{1}$ by several orders of magnitude. Although important, this requirement has been ignored in previous NMR work with quadrupolar nuclei in superconducting compounds.[5, 6]
In Fig. 3(a) the normal state $T_{2}^{-1}$ is field independent, consistent with expectation of magnetic-field-independence of the dipole-dipole interaction as well as of the density of electronic states that determine $T_{1}^{-1}$. In the superconducting state, there is clearly a jump of $T_{2}^{-1}$ at $T_{c}$, shown in Fig. 3(b). The magnitude of the jump is then measured and plotted in the inset of Fig. 3 (a). The fact that the jump scales with the external field is a clear indication of vortex dynamics, since the vortex density is proportional to the field. In the superconducting state, we find a common linear dependence of $T_{2}^{-1}$ on temperature in Fig. 3(b) which means that it scales with magnetic field being plotted versus $T/T_{c}$(H) (Supplementary Materials). Consequently, we also attribute this field dependence to vortex dynamics.
Figure 4: Temperature dependence of 93Nb transverse relaxation rate $T_{2}^{-1}$ of the Nb3Sn powder at 9.43 T from Sample 2. Data are shown as dots with error bars less than dot size. The green and red data are the Lorentzian and Gaussian components of $T_{2}^{-1}$ respectively, achieved from fitting the recovery curve with Eq. 5. The Lorentzian component is further separated into a Redfield component (purple curve) and a dipolar component (blue curve). The Redfield component was calculated from the measurement of $T_{1i}$ plotted in the inset.
With the larger Sample 2 we investigated $T_{2}^{-1}$ behavior in more detail, measuring both $T_{1i}^{-1}$, inset to Fig. 4, and $T_{2}^{-1}$ at 9.43 T with a significant enhancement in the signal-to-noise ratio of the NMR signal, M(t) allowing separation of the $T_{2}^{-1}$ relaxation into two components from the combined function [6],
$M(t)=M_{0}e^{-\frac{t}{T_{2e}}}e^{-(\frac{t}{T_{2g}})^{2}}$ (5)
where $T_{2e}$ (green data) and $T_{2g}$ (blue data) are the Lorentzian and Gaussian components of $T_{2}^{-1}$ respectively. For these data, the Lorentzian Redfield component of $T_{2meas}^{-1}$ was determined from $T_{1i}^{-1}$ and Eq. 4 with $\kappa=0.8$ (purple curve). The latter reflects the constraint of temperature independence of $T_{2}^{-1}$ in the normal state, already evident in the extrapolations in Fig. 3(b). Consequently, the Gaussian dipolar contribution to $T_{2}^{-1}$ is the red data in Fig. 4. After subtraction of the Redfield term, the Lorentzian dipolar contribution is given by the blue curve.
In the normal state, the Gaussian component of $T_{2}^{-1}$ is approximately temperature independent $\sim 1$ ms-1, and much smaller than expected based on the square root of the second moment of the classical dipolar field, at a Nb nucleus in Nb3Sn (indicated as $j$), with the following equation [17]:
$\overline{\Delta\omega^{2}}=M_{2}=(\frac{\mu_{0}}{4\pi})^{2}\frac{3}{4}\gamma^{4}\hbar^{2}I(I+1)\sum_{k}\frac{(1-3\cos^{2}\theta)^{2}}{r_{jk}^{6}},$ (6)
where $\gamma$ is the gyromagnetic ratio, ${\bf r}_{jk}$ is the separation vector of two nuclei, $j$ and $k$, at an angle $\theta$ to the magnetic field, summed over all nuclei $k$ in the Nb3Sn structure. From the calculation, we find $M_{2}^{1/2}=20\,$ms-1, more than an order of magnitude larger than the experimental data shown in Fig. 3 and Fig. 4. This is evidence of a very strong indirect exchange interaction between Nb nuclei [9] reducing the dipolar $T_{2}^{-1}$, and modifying the spectral NMR lineshape to be a mixture of Gaussian and Lorentzian character. [18]
In the superconducting state, the Redfield component is gapped out in the low temperature limit following the behavior of $T_{1i}$, measured independently and shown in the inset to Fig. 4. The dipolar component remains non-zero in the low temperature limit as expected, consistent with the low temperature limit of Sample 1. The Gaussian component decreases toward zero as the temperature decreases, and the Lorentzian component dominates. We attribute this to a modification of the indirect exchange interaction in the superconducting state. This new result may be important for understanding the nature of competition between superconductivity and magnetic RKKY interactions in metallic systems.
Comparing the $T_{2}^{-1}$ result between Sample 1 and Sample 2, we found that Sample 2 does not exhibit a jump at $T_{c}$. Furthermore, the temperature dependence in the superconducting state also appears to be missing a contribution from vortex dynamics. This suggests that vortices are pinned in Sample 2 in contrast to Sample 1, at least on the kHz time scale. From the sample parameters, we know that Sample 2 has a higher critical field and critical temperature (Supplementary Materials) and is closer to being stoichiometric. Comparisons with other superconducting properties of the two samples is beyond the scope of this work.
V Conclusion
In summary, the superconducting order parameter amplitude is strongly suppressed by magnetic field in Nb3Sn as determined from NMR longitudinal relaxation measurements. From transverse relaxation we found an anomalously high nuclear spin coherence, which we attribute to a strong indirect RKKY exchange interaction between Nb nuclei. We discovered that this interaction is modified in the superconducting state. We also identified vortex dynamics and vortex pinning that is sample dependent.
VI Acknowledgement
We thank Jim Sauls for helpful discussions and Michael Brown for providing Sample 1. We also acknowledge assistance from John Scott and Daehan Park in the ultra-low temperature laboratory. This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. The research work at Florida State University was primarily funded by the US. Department of Energy, Office of High Energy Physics, DE-SC0012083.